To get started, Econ Engineering has just borrowed $500,000 which will be paid off in 20 end-of-quarter payments. If the interest is 18% compounded monthly, what will be the size of each loan payment?

Select one:

A. $34,926

B. $38,663

C. $36,833

D. $41,628
Answer #1

ANSWER:

I = 18% PER YEAR OR 18% / 12 = 1.5% PER MONTH AND N = 3 (1 QUARTER HAS 3 MONTHS)

EFFECTIVE RATE = (1 + I) ^ N - 1 = (1 + 1.5%) ^ 3 - 1 = (1.015) ^ 3 - 1 = 1.0456 - 1 = 0.0456 OR 4.56%

LOAN PAYMENT = BORROWED MONEY(A/P,I,N) = 500,000(A/P,4.56%,20) = 500,000 * 0.0773 = 38,663

SO THE LOAN PAYMENT IS $38,663 THAT IS OPTION B.